OpenStudy (anonymous):

solve: 5 over 2x+1 over 4x= 5 over 4+x

OpenStudy (anonymous):

is it ????? 5 5 ---------- = --------- 2x + 1 4 + x ------ 4x

OpenStudy (anonymous):

its 5 1 5 __ x + __ x = __ + x 2 4 4

OpenStudy (anonymous):

Your question can also be written as 5x 1x 5 _____ + ______ = _____ + x 2 4 4 Taking LCD on both sides we get it as 4 for both sides, we get 10x + 1x 5 + 4x 11x 5 + 4x __________ = ---------- ==> ------ = -------- 4 4 4 4 Now multiplying both sides by 4, we cancel out the 4 in the denominators to get 11x = 5 + 4x 11x - 4x = 5 7x = 5 x = 5/7

OpenStudy (anonymous):

if u r satisfied with the answer, pls click GOOD ANSWER button ....☺

jhonyy9 (jhonyy9):

some persons forget making this last step what you have wrote ,,Harkirat" --- why ???